# Policy Mirror Descent with Lookahead

### v1

@@ Appendix C Additional Details and Proofs for Section 6: $h$-PMD with Function Approximation
To incorporate function approximation to $h$-PMD, the overall approach is as follows: compute an estimate of the lookahead value function by fitting a linear estimator using least squares, then perform an update step as in inexact PMD, taking advantage of the same error propagation analysis which is agnostic to the source of errors in the value function. However a problem that arises is in storing the new policy, since a classic tabular storage method would require $\mathcal{O}(|\mathcal{S}|)$ memory, which we would like to avoid to scale to large state spaces. A solution to this problem will be presented later on in this section.
@@ C.1 Inexact $h$-PMD with Linear Function Approximation
Here we provide a detailed description of the $h$-PMD algorithm using function approximation to estimate the $Q_{h}$-function. Firstly we consider the lookahead depth $h$ as a fixed input to the algorithm. We assume that the underlying MDP has finite state and action spaces $\mathcal{S}$ and $\mathcal{A}$, with transition kernel $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{A})$, reward vector $R\in\left[0,1\right]^{|\mathcal{S}|\times|\mathcal{A}|}$ and discount factor $\gamma$. We also assume we are given a feature map $\psi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$, also written as $\Psi\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times d}$.
@@
Computing the full policy $\pi_{k}$ for every state at every iteration would require $\Omega(|\mathcal{S}|)$ operations, although the number of samples to the environment still depends only on $d$. If this is not a limitation the policy update step can be performed as in inexact $h$-PMD simply by using $\Psi\theta_{k}$ as an estimate for $Q_{h}^{\pi_{k}}$, as described in step 3c. Otherwise, we describe a method below to compute $\pi_{k}$ on demand when it is needed in the computation of the targets in step 3a, eliminating step 3c entirely.
@@
This procedure can be used to compute ${\pi_{k+1}}(\cdot|s)$ “on demand” for any $k>0$ and $s\in\mathcal{S}$ by recursively computing ${\pi_{k}}(\cdot|s^{\prime})$ at any states $s^{\prime}$ required for the computation of $\Psi\theta_{k}$ (since $\pi_{0}$ is given). The memory cost requirement for this method is of the order of $\mathcal{O}(K|\mathcal{A}|\min\{|\mathcal{S}_{0}|H,|\mathcal{S}|\})$ in total which is always less than computing the full tabular policy at each iteration.
@@
Above, $(\Psi\theta_{k})_{s}$ denotes a vector in $\mathbb{R}^{|\mathcal{A}|}$ that can be computed in $\mathcal{O}(|\mathcal{A}|d)$ operations: each component in $a$ is equal to $\psi(s,a)^{\top}\theta_{k}$ which can be computed in $\mathcal{O}(d)$ operations. Also, the stepsize $\eta_{k}$ is chosen to be equal to its lower bound in the analysis of 5.3, with the value of $c_{k}$ chosen to be equal to $\gamma^{2h(k+1)}$, as it is needed for linear convergence. We also note that by memorizing policies as they are computed, we manage to perform just as many policy updates as are needed by the algorithm.
C.2 Convergence Analysis of Inexact $h$-PMD using Function Approximation
@@
We have now fully described how to estimate the lookahead value function $Q^{\pi}_{h}$ using a number of trajectories that does not depend on the size of the state space. In the previous section, we described how to use this estimate in an $h$-PMD style update, while maintaining a memory size that does not depend on $|\mathcal{S}|$. Therefore we have an $h$-PMD algorithm that does not use memory, computation or number of samples depending on $|\mathcal{S}|$.
Appendix D Additional Details and Simulations
Here we provide details of an additional experiment running $h$-PMD with Euclidean Divergence. As in section 7, experiments were run on the “bsuite” DeepSea environment with $\gamma=0.99$ and a grid size of 64x64. The figures are identical in structure to the ones in 7, 32 runs were performed for each value of $h$ in 1, 5, 10, 15, and 20. In Fig. 3, we report the number of samples to the environment at each iteration for the same runs. When compared to the figures showing the convergence, it is clearly seen that running the algorithm with higher values of lookahead results in convergence using fewer overall samples.

### v2

@@ Appendix C Additional Details and Proofs for Section 6: $h$-PMD with Function Approximation
To incorporate function approximation to $h$-PMD, the overall approach is as follows: compute an estimate of the lookahead value function by fitting a linear estimator using least squares, then perform an update step as in inexact PMD, taking advantage of the same error propagation analysis which is agnostic to the source of errors in the value function. However a problem that arises is in storing the new policy, since a classic tabular storage method would require $\mathcal{O}(S)$ memory, which we would like to avoid to scale to large state spaces. A solution to this problem will be presented later on in this section.
@@ C.1 Inexact $h$-PMD with Linear Function Approximation
Here we provide a detailed description of the $h$-PMD algorithm using function approximation to estimate the $Q_{h}$-function. Firstly we consider the lookahead depth $h$ as a fixed input to the algorithm. We assume that the underlying MDP has finite state and action spaces $\mathcal{S}$ and $\mathcal{A}$, with transition kernel $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{A})$, reward vector $R\in\left[0,1\right]^{S\times A}$ and discount factor $\gamma$. We also assume we are given a feature map $\psi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$, also written as $\Psi\in\mathbb{R}^{SA\times d}$.
@@
Computing the full policy $\pi_{k}$ for every state at every iteration would require $\Omega(S)$ operations, although the number of samples to the environment still depends only on $d$. If this is not a limitation the policy update step can be performed as in inexact $h$-PMD simply by using $\Psi\theta_{k}$ as an estimate for $Q_{h}^{\pi_{k}}$, as described in step 3c. Otherwise, we describe a method below to compute $\pi_{k}$ on demand when it is needed in the computation of the targets in step 3a, eliminating step 3c entirely.
@@
This procedure can be used to compute ${\pi_{k+1}}(\cdot|s)$ “on demand” for any $k>0$ and $s\in\mathcal{S}$ by recursively computing ${\pi_{k}}(\cdot|s^{\prime})$ at any states $s^{\prime}$ required for the computation of $\Psi\theta_{k}$ (since $\pi_{0}$ is given). The memory cost requirement for this method is of the order of $\mathcal{O}(KA\min\{|\mathcal{S}_{0}|H,S\})$ in total which is always less than computing the full tabular policy at each iteration.
@@
Above, $(\Psi\theta_{k})_{s}$ denotes a vector in $\mathbb{R}^{A}$ that can be computed in $\mathcal{O}(Ad)$ operations: each component in $a$ is equal to $\psi(s,a)^{\top}\theta_{k}$ which can be computed in $\mathcal{O}(d)$ operations. Also, the stepsize $\eta_{k}$ is chosen to be equal to its lower bound in the analysis of 5.3, with the value of $c_{k}$ chosen to be equal to $\gamma^{2h(k+1)}$, as it is needed for linear convergence. We also note that by memorizing policies as they are computed, we manage to perform just as many policy updates as are needed by the algorithm.
C.2 Discussion of Assumption 6.2
We make two preliminary comments before discussing the relevance of our assumption and alternatives to relax it below:
  • First, notice that we do not require $Q^{\pi}_{h}$ to be represented as a linear function. Our assumption is an approximate action value function approximation.
  • When the lookahead depth is $h=1$, we recover the standard PMD algorithm. Notice that in this case our assumption is standard in the RL literature for linear function approximation.
Now, using our notations, notice that for any policy $\pi\in\Pi$, $Q_{h}^{\pi}=r+\gamma P\mathcal{T}^{h-1}V^{\pi}=r+\gamma P\mathcal{T}^{h-1}M^{\pi}Q^{\pi}=J\mathcal{T}^{h-1}M^{\pi}Q^{\pi}.$ We have two alternatives:
  • Directly approximate $Q_{h}^{\pi}$ as we do in our assumption. In that case the targets used in our regression procedure are estimates $\hat{Q}_{h}$ of $Q_{h}^{\pi}$. Note also that this approach is meaningful from a practical point as we may directly use a neural network to approximate the lookahead Q-function which is used in our $h$-PMD algorithm rather than approximating the Q-function first and then trying to approximate from it the lookahead Q-function $Q_{h}^{\pi}.$
  • Approximate $Q^{\pi}$ with linear function approximation and use our propagation analysis (see section 5.1 and proof in appendix B.3.1, B.3.2) based on the link between $Q_{h}^{\pi}$ and $Q^{\pi}$ highlighted in the formula above. This approach which propagates the function approximation error on $Q^{\pi}$ to $Q_{h}^{\pi}$ constitutes an alternative to Assumption 6.2 on $Q_{h}^{\pi}$ by an assumption on $Q^{\pi}.$ We discuss it further in the rest of this section.
In the following exposition we provide more details regarding estimating $Q^{\pi}$ using linear function approximation and relaxing assumption 6.2 using the second approach discussed above.
Let Assumption 6.1 hold (as in Theorem 6.3). Instead of assumption 6.2, we introduce the following assumption on the $Q$-function $Q^{\pi}$ rather than the $h$-step lookahead Q-function $Q_{h}^{\pi}\,.$
Assumption C.3.
$\exists\,\epsilon>0,\forall\pi\in\Pi,\inf_{\theta\in\mathbb{R}^{d}}\lVert Q^{\pi}-\Psi\theta\rVert\leq\epsilon\,.$
We modify the $h$-PMD with function approximation algorithm as follows: instead of directly approximating $Q_{h}^{\pi}$ using linear function approximation, we first use linear function approximation to approximate $Q^{\pi}$ from which we construct an estimate of $Q_{h}^{\pi}$ similarly to our method for inexact $h$-PMD. This requires us to replace the regression targets with simpler ones, which are Monte Carlo estimates of $Q^{\pi}$ instead of $Q^{\pi}_{h}$. Specifically, these new targets are defined as:
$R_{M_{0}}(z):=\frac{1}{M_{0}}\sum_{i=1}^{M_{0}}\sum_{t=0}^{H-1}\gamma^{t}r(s_{t}^{(i)},a_{t}^{(i)}),$
where $(s_{t}^{(i)},a_{t}^{(i)})$ are state action pairs sampled according to the policy $\pi$, $z=(s,a)\in\mathcal{C}\subseteq\mathcal{S}\times\mathcal{A}$ and $(s_{0}^{(i)},a_{0}^{(i)})=z$.
For any policy $\pi$ we can use the same function approximation procedure as described in our paper (see Appendix C.1) to yield a $\hat{\theta}\in\mathbb{R}^{d}$ with $\lVert Q^{\pi}-\Psi\hat{\theta}\rVert_{\infty}\leq\epsilon_{0}$, (for some value of $\epsilon_{0}$ to be defined below) by simply plugging in $h=1$. Note that this particular case was previously analysed in Szepesvári [2023] (Lecture 8). We reuse an intermediate result therein (equation (7)) which guarantees the bound:
$\lVert Q^{\pi}-\Psi\hat{\theta}\rVert_{\infty}\leq\lVert\epsilon_{\pi}\rVert_{\infty}(1+\sqrt{d})+\sqrt{d}\left(\frac{\gamma^{H}}{1-\gamma}+\frac{1}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{C}|/\delta)}{2M_{0}}}\right)=:\epsilon_{0}(\delta)$ (59)
with probability at least $1-\delta$ for some $10<\delta<$ 1.
Now we reuse most of our estimation procedure in Section 5.1, simply by changing the definition of $\hat{V_{1}}$ to $\hat{V_{1}}(s):=\sum_{a\in\mathcal{A}}\pi(a|s)(\Psi\hat{\theta})_{s,a}$ for all $s\in\mathcal{S}_{0}$. This evidently retains the guarantee that $|\hat{V}_{1}(s)-V^{\pi}(s)|\leq\epsilon_{0}(\delta)$ with probability at least $1-\delta$ for each $s\in\mathcal{S}_{0}$. Finally, using the same error propagation analysis as in Appendix B.3.1, replacing $C_{1}^{V}(\frac{\delta_{1}^{V}}{|\mathcal{S}_{0}|})$ with $\epsilon_{0}(\delta_{1}^{V})$ we obtain the following bound for any individual state action pair $(s,a)$:
$|\hat{Q^{\pi}_{h}}(s,a)-Q^{\pi}_{h}(s,a)|\leq\frac{\gamma}{1-\gamma}\sqrt{\frac{\log(2/\delta_{h}^{Q})}{M_{h}}}+\gamma^{2}\frac{1-\gamma^{h-1}}{(1-\gamma)^{2}}\sqrt{\frac{\log(2(h-1)Z/\delta_{t}^{V})}{M}}+\gamma^{h}\epsilon_{0}(\delta_{1}^{V})$ (60)
with probability at least $1-\delta_{1}^{V}-\delta_{t}^{V}-\delta_{h}^{Q}$.
Finally, the update rule as described in Appendix C.1 is replaced by $\pi_{k+1}(\cdot|s)\in\operatorname{argmax}_{\pi_{s}\in\Delta(\mathcal{A})}(\eta_{k}\langle\hat{Q_{h}^{\pi_{k}}}(s,\cdot),\pi_{s}\rangle-D_{\phi}(\pi_{s},\pi_{s}^{k}))$. Note that we only compute $\pi_{k}$ in the states where we need it to compute $\pi_{k+1}$ as described in Appendix C.1.
Unfortunately, it is not clear how to directly obtain a result similar to Theorem 6.3 from this approach (controlling $\lVert V^{\star}-V^{\pi_{k}}\rVert_{\infty}$ at each iteration, i.e. uniformly over all states) beyond the bound that we provide above. A naive union bound over the state action space would result in a bound with a logarithmic dependence on the size of the state space, unlike Theorem 6.3 which results in a bound that is completely independent of the size of the state space.
We would like to add an additional comment regarding possible ways to relax Assumption 6.2 even for the standard case where $h=1$. Recently, Mei et al. Mei et al. [2024] proposed an interesting alternative approach relying on order-based conditions (instead of approximation error based ones) to analyze policy gradient methods with linear function approximation. We would like to highlight though that their work only applies to the bandit setting to the best of our knowledge. The extension to MDPs is far from being straightforward and would require some specific investigation. It would be interesting to see if one can relax Assumption 6.2 using similar ideas.
C.3 Convergence Analysis of Inexact $h$-PMD using Function Approximation
@@
We have now fully described how to estimate the lookahead value function $Q^{\pi}_{h}$ using a number of trajectories that does not depend on the size of the state space. In the previous section, we described how to use this estimate in an $h$-PMD style update, while maintaining a memory size that does not depend on $S$. Therefore we have an $h$-PMD algorithm that does not use memory, computation or number of samples depending on $S$.
Appendix D Further Simulations
D.1 Simulations for $h$-PMD with Euclidean Regularization
Here we provide details of an additional experiment running $h$-PMD with Euclidean regularization. As in section 7, experiments were run on the bsuite DeepSea environment Osband et al. [2020] with $\gamma=0.99$ and a grid size of 64x64. The figures are identical in structure to the ones in 7, 32 runs were performed for each value of $h$ in 1, 5, 10, 15, and 20. In Fig. 3, we report the number of samples to the environment at each iteration for the same runs. When compared to the figures showing the convergence, it is clearly seen that running the algorithm with higher values of lookahead results in convergence using fewer overall samples.
@@ D.1 Simulations for $h$-PMD with Euclidean Regularization
D.2 Sample Complexity Comparison for Different Lookahead Depths
We have recorded the sample complexity of our algorithm for different values of $h$ in Figure 3. Notice that the number of samples used at each iteration increases with $h$. However, notice that this increase is small: $h$-PMD with $h=20$ requires only 1.5 times more samples per iteration than $h=1$ as shown in Figure 3. Furthermore, increasing $h$ greatly improves the convergence rate of the algorithm as described by our theory, which results in a much smaller number of iterations required until convergence (for example $h=20$ converges in at least 10 times fewer iterations than $h=1$ (see figure 1 (right) in section 7). Overall, the algorithm is more sample efficient for higher values of $h$.
@@ D.2 Sample Complexity Comparison for Different Lookahead Depths
D.3 Total Running Time Comparison for Different Lookahead Depths
Figure 4: Value function gap against running time. Note that the algorithms with higher values for $h$ still converge faster in terms of runtime.
We provide a plot accounting for running time instead of the number of iterations (see Figure 4). Observe that the overall performance is very similar to the same plot with the number of iterations on the x-axis.
D.4 About the Stepsizes in Simulations
In our experiments we use the exact same stepsizes described in our theory. We also observe in our simulations that the stepsize does not go to infinity but rather decreases at some point and/or stabilizes. This is a consequence of the adaptivity of the stepsize discussed above. We also note that the simulations reported in Johnson et al. [2023] (appendix D p. 18) also validate the fact that the stepsizes do not go to infinity (see the right plots therein for the stepsizes). We also observe a similar behavior for the stepsizes in our experiments (see Figure 5).
Figure 5: Behaviour of our adaptive stepsize $\eta_{k}$ at each iteration of the algorithm for different values of $h$. Note that the stepsize does not diverge towards infinity, instead it usually vanishes after a certain number of iterations.
D.5 Larger lookahead depths
In order to practically investigate the tradeoff of increasing the lookahead depth $h$, we performed additional experiments where $h$ was increased to $100$ in the DeepSea bsuite environment. The plots in Figure 6 show that increasing $h$ past the values in our initial experiments still results in better performance.
Figure 6: Value function gap plotted against number of iterations (left) and against running time (right) of $h$-PMD in the DeepSea bsuite environment. Note that increasing the lookahead depth seems to result in better performance in terms of runtime and number of iterations.
D.6 Experiments with MCTS-based Lookahead Estimation
We have performed additional experiments in an even more challenging and practical setting, on two environments from the bsuite Osband et al. [2020] set: DeepSea and Catch. We have implemented our PMD algorithm with a Monte Carlo Tree Search (MCTS) algorithm (which we call PMD-MCTS) to compute the lookahead Q-function. We used DeepMind’s MCTS implementation (MCTX) [^3] https://github.com/google-deepmind/mctx in order to run this algorithm on any deterministic gym style environments implemented in JAX (we use gymnax Lange [2022] to implement these environments). Note that any tree search algorithm implementation can be used to estimate the lookahead function and plugged into our PMD method. See figures below. Note that this training procedure is more practical: we do not evaluate the value function at each state-action pair anymore at each time step, we only evaluate the value function (and so only update the policy) in states that we encounter using the current policy. This significantly relaxes the generative model assumption. Note that in this more challenging setting the best performance is not obtained for higher values of $h$: intermediate values of $h$ perform better, illustrating the tradeoff in choosing the depth $h$.
Figure 7: Training curves of PMD-MCTS for the bsuite Catch (left) and DeepSea (right) environments. Note that the case of $h=1$ fails to converge in any reasonable number of iterations in both environments.
D.7 Continuous Control Experiments
Finally, we performed experiments in classic continuous control environments in order to investigate the performance of $h$-PMD when applied to MDPs with large/infinite state spaces. For this continuous state space setting, we implemented a fully online variant of $h$-PMD based off of DeepMind’s MuZero Schrittwieser et al. [2019] algorithm. Experiments show that increasing the lookahead depth can lead to better performance up to a certain point, and the tradeoff seems to be environment-specific.
Figure 8: Performance of $h$-PMD in two of OpenAI’s gym Brockman et al. [2016] environments, CartPole-v1 (left) and Acrobot-v1 (right). In both figures the reward is plotted against the total number of environment steps evaluated, which can be understood as the number of samples used. Note that higher values of $h$ do lead to better performance for the CartPole-v1 environment, but not necessarily for the Acrobot-v1 environment.
D.8 Additional experiments
We conclude with results from a larger set of experiments designed to verify the robustness of our experimental results to different environment parameters. We report the performance of the $h$-PMD algorithm in the inexact setting in two different environments, DeepSea from bsuite Osband et al. [2020], and the classic Gridworld environment. The effect of $\gamma$ is as expected: higher values of $\gamma$ result in slower convergence, lower values of $\gamma$ yield faster convergence. The size of the environment (“chain length” for DeepSea and number of columns/rows for Gridworld) directly affects the number of iterations until convergence predictably: larger environments take more iterations to solve. Note that in all cases the effect of $h$ is as expected, i.e. higher $h$ leads to faster convergence.